# EGA: A Unified End-to-End Generative Framework for Industrial Advertising Systems

### v1

@@ Abstract.
Online industrial advertising system is fundamentally constrained by the inefficiency of multi-stage cascaded architectures, which filter out high-potential candidates early and fragment business decision logic across independent modules. Although recent advances in generative recommendation offer end-to-end solutions, they fall short of practical advertising requirements, lacking explicit modeling of bidding, creative selection, allocation mechanism, and payment computation that are essential for real-world deployment.
To overcome these limitations, we propose End-to-end Generative Advertising (EGA), a first unified generative framework that seamlessly integrates user interests modeling, POI and creative generation, position allocation, and payment optimization within a single model. EGA leverages hierarchical tokenization and multi-token prediction to jointly generate candidate POI and creative contents, while a permutation-aware reward model and token-level bidding strategy ensure alignment with both user experiences and advertiser business objectives. Meanwhile, we decouple allocation from payment via a dedicated POI-level payment network with differentiable ex-post regret minimization, guaranteeing incentive compatibility approximately. Extensive offline and large-scale online experiments on real-world advertising systems demonstrate EGA’s effectiveness and practical advantages over traditional cascading architectures, highlighting its potential as one of the industry’s pioneering end-to-end generative advertising solutions.
@@ 5. Experiments
  • RQ3: How is the scalability of the EGA model?
@@ 5.4. Model Scalability (RQ3)
We conduct a comprehensive study to evaluate the sensitivity and scalability of EGA with respect to key hyperparameters and model size. The results are summarized in Figure 5 and Figure 6, where RPM and CTR-poi are used as the evaluation metrics. Figure 5 examines the impact of token-level bidding parameters $\alpha$ and $\beta$ in Equation (21). $\alpha$ measures the importance of bids during the allocation process. Increasing $\alpha$ significantly boosts RPM at first, as higher-bid ads become more likely to win in the generative allocation. However, this trend gradually plateaus as further increasing $\alpha$ provides diminishing marginal returns. In contrast, CTR-poi shows a consistent decline as $\alpha$ rises, reflecting a shift in exposure from more organic contents to higher-bid ads. Besides, $\beta$ acts as a global balancing factor between ads and organic content. Increasing $\beta$ reduces the probability of ads being allocated. As a result, RPM decreases steadily with increasing $\beta$, and eventually flattens out as ads rarely win positions in the list. Conversely, CTR-poi increases with $\beta$, as the display list becomes dominated by organic content. In practice, the optimal values of $\alpha$ and $\beta$ are selected to achieve a trade-off between platform revenue and user experience.
Figure 6. Scalability of EGA on model scaling.
Figure 6 explores the effect of model size. Increasing the number of encoder/decoder blocks leads to a notable improvement in RPM in the early stage, followed by a plateau as the model becomes deeper, reflecting the law of diminishing returns. Similarly, expanding the length of user behaviors results in a rapid rise in RPM at first, then levels off as more behaviors are added. This demonstrates that while both model capacity and data scale are crucial for performance, their incremental benefit reduces as they become sufficiently large.
@@ 6. Conclusion
In this work, we presented End-to-end Generative Advertising (EGA), a novel framework that unifies ranking, creative selection, ad allocation, and payment computation into a single generative model for industrial advertising systems. By leveraging hierarchical semantic tokenization, permutation-aware reward modeling, and token-level bidding and allocation, EGA bridges the gap between user interests modeling and business-critical auction constraints such as IC and IR. The proposed multi-phase training paradigm, including interest-based pre-training and auction-based post-training, ensures that EGA captures both user interests and advertiser utility under complex real-world conditions. Extensive offline experiments and large-scale online A/B tests on an industrial dataset demonstrate that EGA achieves substantial improvements over both multi-stage cascading architecture and recent generative recommendation baselines in platform revenue, user experience, and advertiser return on investment, all with minimal latency overhead. Our ablation studies validate the effectiveness of each design component and highlight the flexibility and scalability of the proposed framework.

### v2

# Beyond Cascaded Architectures: An End-to-end Generative Framework for Industrial Advertising

@@ Abstract.
Traditional online industrial advertising systems suffer from the limitations of multi-stage cascaded architectures, which often discard high-potential candidates prematurely and distribute decision logic across disconnected modules. While recent generative recommendation approaches provide end-to-end solutions, they fail to address critical advertising requirements of key components for real-world deployment, such as explicit bidding, creative selection, ad allocation, and payment computation.
To bridge this gap, we introduce End-to-End Generative Advertising (EGA), the first unified framework that holistically models user interests, point-of-interest (POI) and creative generation, ad allocation, and payment optimization within a single generative model. Our approach employs hierarchical tokenization and multi-token prediction to jointly generate POI recommendations and ad creatives, while a permutation-aware reward model and token-level bidding strategy ensure alignment with both user experiences and advertiser objectives. Additionally, we decouple allocation from payment using a differentiable ex-post regret minimization mechanism, guaranteeing approximate incentive compatibility at the POI level. Through extensive offline evaluations and large-scale online experiments on real-world advertising platforms, we demonstrate that EGA significantly outperforms traditional cascaded systems in both performance and practicality. Our results highlight its potential as a pioneering fully generative advertising solution, paving the way for next-generation industrial ad systems.
@@ 5. Experiments
  • RQ3: How do hyperparameters affect model performance?
@@ 5.4. Hyperparameters (RQ3)
We conduct a comprehensive study to evaluate the sensitivity of EGA with respect to key hyperparameters. The results are summarized in Figure 5, where RPM and CTR-poi are used as the evaluation metrics. Figure 5 examines the impact of token-level bidding parameters $\alpha$ and $\beta$ in Equation (21). $\alpha$ measures the importance of bids during the allocation process. Increasing $\alpha$ significantly boosts RPM at first, as higher-bid ads become more likely to win in the generative allocation. However, this trend gradually plateaus as further increasing $\alpha$ provides diminishing marginal returns. In contrast, CTR-poi shows a consistent decline as $\alpha$ rises, reflecting a shift in exposure from more organic contents to higher-bid ads. Besides, $\beta$ acts as a global balancing factor between ads and organic content. Increasing $\beta$ reduces the probability of ads being allocated. As a result, RPM decreases steadily with increasing $\beta$, and eventually flattens out as ads rarely win positions in the list. Conversely, CTR-poi increases with $\beta$, as the display list becomes dominated by organic content. In practice, the optimal values of $\alpha$ and $\beta$ are selected to achieve a trade-off between platform revenue and user experience.
@@ 6. Conclusion
In this work, we presented End-to-end Generative Advertising (EGA), a novel framework that unifies ranking, creative selection, ad allocation, and payment computation into a single generative model for industrial advertising systems. By leveraging hierarchical semantic tokenization, permutation-aware reward modeling, and token-level bidding and allocation, EGA bridges the gap between user interests modeling and business-critical auction constraints such as IC and IR. The proposed multi-phase training paradigm, including interest-based pre-training and auction-based post-training, ensures that EGA captures both user interests and advertiser utility under complex real-world conditions. Extensive offline experiments and large-scale online A/B tests on an industrial dataset demonstrate that EGA achieves substantial improvements over both multi-stage cascading architecture and recent generative recommendation baselines in platform revenue, user experience, and advertiser return on investment, all with minimal latency overhead. Our ablation studies validate the effectiveness of each design component and highlight the flexibility of the proposed framework.